OpenStudy (chycora):

Factor. 6x² - 46x - 16 A. 2(x - 8)(3x + 1) B. 2(x + 4)(3x - 4) C. 2(x - 8)(3x - 1) D. 2(x + 1)(3x - 8)

OpenStudy (chycora):

@Ajohnson2199

OpenStudy (anonymous):

what do you think

OpenStudy (anonymous):

if you don't know I can help

OpenStudy (chycora):

i think it's a or c

OpenStudy (anonymous):

6x² - 46x - 16 = 2(3x² - 23x - 8) = 2(3x² - 24x + x - 8) = 2[3x(x - 8) + (x - 8) = 2(x - 8)(3x + 1) Answer choice: A.2(x - 8)(3x + 1) yes A is the answer

OpenStudy (anonymous):

medal please

OpenStudy (anonymous):

thanks and I gave you one

OpenStudy (anonymous):

you guessed correctly so you got a medal

OpenStudy (chycora):

well thanx